Teachers’Salaries The average annual salary for all U.S. teachers is $47,750. Assume that the distribution is normal and the standard deviation is $5680. Find the probability that a randomly selected teacher earns


a. Between $35,000 and $45,000 a year


b. More than $40,000 a year


c. If you were applying for a teaching position and were offered $31,000 a year, how would you feel

1
Expert's answer
2022-01-13T16:24:17-0500

a. "P(35000<X<45000)=P(\frac{35000-47750}{5680}<Z<\frac{45000-47750}{5680})="

"=P(-2.24<Z<-0.48)=P(Z<-0.48)-P(Z<-2.24)="

"=P(Z<2.24)-P(Z<0.48)=0.3031."


b. "P(X>40000)=P(Z>\frac{40000-47750}{5680})=P(Z>-1.36)="

"=P(Z<1.36)=0.9131."


c. "P(X<31000)=P(Z<\frac{31000-47750}{5680})=P(Z<-2.95)="

"=1-P(Z<2.95)=0.0016."

Since this probability is very unusual, you would feel very bad.